OpenStudy (anonymous):

Can anyone help me solve this equation -6+3x=-9 A)-1 B)-6 C)-5 D)-3

OpenStudy (anonymous):

Main Objective: Get the unknown variable by itself

OpenStudy (anonymous):

To do this we want to add or subtract the constant from both sides

OpenStudy (anonymous):

-6 + 3x = -9 Add 6 to each side. 3x = -3 Divide 3 from each side. x = -1 So, A) -1 is the correct answer.

OpenStudy (anonymous):

arrange the equation so that x is on on side and the numbers on the other side

OpenStudy (anonymous):

This will eliminate the constant from the left side of the equation. Then we can divide both sides by the coefficient to get the value of x

OpenStudy (anonymous):

a

OpenStudy (anonymous):

-6+3x=-9 or 3x=-9+6 3x=-3 x=-1

OpenStudy (anonymous):

Thnks so, so much u guys r great :-D